use method of substitution to solve the system of linear equations.
7X+5Y=7/2
Y=2X-1

Answer by flame8855(424)

y = 2x-1
7x+5y=7/2
7x+5(2x-1) = 7/2
17x= 17/2
x= 1/2
y=0